8. A certain pure solvent freezes at 39.8°C and has a freezing point depression constant Kp = 0.777°C/m. What is the predicted fPlease help is due in 30 mins, will rate, Thank you!
Answer #1

Answer:

Given that t

freezing point depression constant, Kf = 0.777 C/m

Normal freezing point = 39.8 degree centigrade

Concentration in molality = 1.42 × 100 = 1.42 m

So, (delta T) = Kf × m = 0.777 × 1.42 = 1.103

So, the freezing point is lowered by 1.103 degree centigrade.

So the predicted freezing point is = 39.8 - 1.103

= 38.69 degree centigrade.